OpenStudy (anonymous):

Multiplying polynomials: (x^2+3)(x-4) I have a few more problems after this one....

OpenStudy (anonymous):

Thanks @bohotness

OpenStudy (anonymous):

Im still kind of confused on how I would combine them if they are not like terms. Like could I multiply x by y?

OpenStudy (academicgurusinc):

In this case, you have to take everything step-by-step. So you'll multiply your first term "X" to the other parenthesis. You multiply it to every item in the second parenthesis. So you'll have (x^2+3)(x-4) = x^3-4x Now, you'll multiply your second term "3" to the second parenthesis. So: (x^2+3)(x-4)=x^3-4x^2+3x-12 is your final answer. Hope that helps! If you found this helpful, I encourage to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)